Hi, I am having troubles with finding "unit sales of 14,000 and 16,000 (flexible budget) for direct materials etc.

Phoenix Company’s 2017 master budget included the following fixed budget report. It is based on an expected production and sales volume of 15,000 units. PHOENIX COMPANY Fixed Budget Report For Year Ended December 31, 2017 Sales $ 3,150,000 Cost of goods sold Direct materials $ 975,000 Direct labor 240,000 Machinery repairs (variable cost) 60,000 Depreciation—Plant equipment (straight-line) 315,000 Utilities ($30,000 is variable) 210,000 Plant management salaries 210,000 2,010,000 Gross profit 1,140,000 Selling expenses Packaging 90,000 Shipping 105,000 Sales salary (fixed annual amount) 235,000 430,000 General and administrative expenses Advertising expense 125,000 Salaries 241,000 Entertainment expense 85,000 451,000 Income from operations $ 259,000 Required: 1&2. Prepare flexible budgets for the company at sales volumes of 14,000 and 16,000 units and classify all items listed in the fixed budget as variable or fixed.
Answer #1
PHOENIX COMPANY
Flexible Budget
For Year Ended December 31, 2017
Flexible Budget Flexible Budget for:
Variable Amount per Unit Total Fixed Cost Units Sales of 14,000 Unit Sales of 16,000
Sales 210 2940000 3360000
Variable costs:
Direct materials 65 910000 1040000
Direct labor 16 224000 256000
Machinery repairs 4 56000 64000
Utilities 2 28000 32000
Packaging 6 84000 96000
Shipping 7 98000 112000
Total variable costs 100 1400000 1600000
Contribution margin 110 1540000 1760000
Fixed costs
Depreciation—Plant equipment (straight-line) 315000 315000 315000
Advertising expense 125000 125000 125000
Entertainment expense 85000 85000 85000
Plant management salaries 210000 210000 210000
Utilities 180000 180000 180000
Sales salary 235000 235000 235000
Salaries 241000 241000 241000
Total fixed costs 1391000 1391000 1391000
Income from operations 149000 369000
Workings:
Sales = 3150000/15000 = $210
Variable costs:
Direct materials=975000/15000 = $65
Direct labor=240000/15000 = $16
Machinery repairs=60000/15000= $4
Utilities=30000/15000 = $2
Packaging=90000/15000=$6
Shipping=105000/15000 = $7